# LearnLM: Improving Gemini for Learning

### v2

@@ Abstract
Today’s generative AI systems are tuned to present information by default rather than engage users in service of learning as a human tutor would. To address the wide range of potential education use cases for these systems, we reframe the challenge of injecting pedagogical behavior as one of pedagogical instruction following, where training and evaluation examples include system-level instructions describing the specific pedagogy attributes present or desired in subsequent model turns. This framing avoids committing our models to any particular definition of pedagogy, and instead allows teachers or developers to specify desired model behavior. It also clears a path to improving Gemini models for learning—by enabling the addition of our pedagogical data to post-training mixtures—alongside their rapidly expanding set of capabilities. Both represent important changes from our initial tech report [1]. We show how training with pedagogical instruction following produces a LearnLM model (available on Google AI Studio) that is preferred substantially by expert raters across a diverse set of learning scenarios, with average preference strengths of 31% over GPT-4o, 11% over Claude 3.5, and 13% over the Gemini 1.5 Pro model LearnLM was based on.
@@ 1 Introduction
Figure 1: An overview of our three-stage human evaluation pipeline and our results for comparing LearnLM with other systems. (1) Learning scenarios are developed that allow raters to role-play specific learners interacting with pairs of AI tutors (2). Grounding material (e.g. an essay, homework problem, diagram, etc.) and System Instructions specific to each scenario are passed as context to each model. The resulting conversation pairs are reviewed by pedagogy experts (3) who answer a range of questions assessing each model on its own as well as their comparative performance. These comparative ratings (on a seven-point -3 to +3 Likert scale) are aggregated (4) to show overall preference for LearnLM over GPT-4o, Claude 3.5, and Gemini 1.5 Pro. See Section 4 for more detailed results.
Using the updated methodology, we trained a new version of LearnLM, which is based on Gemini 1.5 Pro [^2] Specifically gemini-1.5-pro-002 (release notes). [3]. In our evaluations against contemporaneous flagship models, each representing a company’s premier offering as of 2024-10-01, expert pedagogical raters preferred LearnLM with an average preference strength of 31% over GPT-4o, 11% over Claude 3.5 Sonnet, and 13% over the original Gemini 1.5 Pro (see Figure 1). LearnLM is available as an experimental model on Google AI Studio along with documentation of example use cases and suggested prompts. We welcome any feedback, which will help inform our future research and prioritization. As we improve LearnLM for teaching and learning, we are also working to bring these advances into Gemini models, so any developers using Gemini can benefit from the improvements made via LearnLM research.
Section 2 describes how we trained LearnLM for pedagogical instruction following and Section 3 explains how we updated our scenario-based evaluation design accordingly. Section 4 shows a detailed analysis of results comparing LearnLM with other premier model offerings. Finally, Section 5 outlines some future work, especially with regards to continued evaluation. Training and evaluation of LearnLM is done across a broad range of core academic subjects, and we include a feasibility study on medical education subjects in Appendix C.
@@ 2.3 Benefits of co-training
3 Human Evaluation Design
In our initial tech report we discussed a taxonomy of pedagogy evaluation designs, and reported results of four human evaluations with different methodologies (Sections 4 and 5 in Jurenka et al. [1]). Here, we focus on scenario-guided, conversation-level pedagogy evaluations and side-by-side comparisons. We improved the clarity and coverage of our learning scenarios, added System Instructions specific to each scenario, and updated the pedagogy rubric and questions. Guiding the conversations with scenarios is especially important in multi-turn settings [10]: without scenarios, the unconstrained nature of human-AI interactions frequently leads to meandering conversations, offering a poor basis for comparison. In contrast, scenario-based approaches support relatively repeatable, controlled comparisons of the capabilities of different conversational AI systems. Scenario frameworks also help with evaluation coverage, ensuring that we test a diverse range of use cases.
Our evaluation process takes place in three stages, depicted above in Figure 1. First, we identified an ecologically representative distribution of learning use cases and created a bank of 49 evaluation scenarios (Section 3.1). Second, these scenarios grounded interactions between AI systems and a pool of $N=186$ pedagogy experts role-playing as learners across learning goals, subjects, learning materials, and learner personas (Section 3.2). Third, to assess the quality of pedagogy in these interactions, we separately recruited a pool of $N=248$ pedagogy experts to review the performance of the systems (Section 3.3). This process produced ample quantitative and qualitative data to help us understand the systems’ capabilities and behavior (Section 3.4).
We are committed to following best practices in research ethics, including by communicating transparently about our research aims, collecting informed consent, and compensating fairly for participation [11]. Our protocol underwent independent ethical review, with a favourable opinion from the Human Behavioural Research Ethics Committee at Google DeepMind (#21 008).
@@ 3.1 Scenario design
We next collaboratively and iteratively populated our bank of scenarios. Members of our team, including two with many years of professional experience in education of students as well as teachers, independently drafted scenarios, leveraging the template and guiding questions from Phase 2. We collectively reviewed the scenario drafts, assessing each for clarity, completeness, correctness, and relevance to our pedagogical principles and the use cases defined in Phase 1. We weighted the overall distribution of scenarios across different learning goals, personas, and subject areas, flagging any gaps for further development. This process resulted in a diverse bank of 49 scenarios across academic subjects (see Appendix B.3 for examples).
@@ 3.2 Conversation collection
Every session of conversation collection began with a short training on role-playing the scenarios (see Figure 2, Step 1). After passing a quiz at the end of the training, participants selected a scenario to enact (see Figure 2, Step 2). Conversation collection proceeded in pairs, such that the same participant enacted a scenario first with one AI system, and then another (LearnLM and a comparison system). We randomized the order of the systems for each conversation pair and did not label the systems for participants. Within each pair of conversations, the models received the same System Instructions, grounding material, and initial learner queries as context, as specified by our scenarios (see Figure 2, Step 3). We formatted all inputs identically, except for some small specification differences mandated by the system APIs.
As specified by our template, each scenario included an initial query for the learner that was automatically sent on behalf of the participant to begin the conversation. After the AI system responded to that query, the participant continued the interaction, guided by the information provided in the scenario. We required participants to continue for a minimum of 10 conversational turns (thus, a minimum of 5 learner and 5 system turns) before they could end the interaction, but on average, participants conversed with the models for around twice this number of turns (Figure 3).
After ending each conversation (see Figure 2, Step 4), participants filled out a brief questionnaire to share their experience interacting with the system (see Figure 2, Step 5 & Appendix B.4). Additionally, after each pair of conversations, participants completed another questionnaire focused on their impressions comparing the two systems (see Figure 2, Step 6 & Appendix B.5). Participants could then either select a new scenario to begin two additional conversations or end the session.
@@ 3.4 Analysis
Our study design involves repeated measurements from our participants. That is, each participant role-playing as a learner interacted with each system multiple times, and each expert assessed each system multiple times. To account for this non-independence and avoid artificially inflating our confidence in our estimates, we analyze our data with hierarchical models [12].
In addition, we conducted qualitative analysis of the open-ended comments and feedback collected from our experts after role-playing each scenario with two systems (stage 2) [^3] We collected open-ended feedback in stage 3 as well, but because the evaluation rubric in this stage is quite extensive, the raters did not provide much additional detail in their comments.. To do so, we first identified and then refined general themes related to the learner-system interactions from participants’ free-form responses. We then coded individual responses for the presence or absence of each theme. To avoid biasing our annotations, we censored the identities of the systems during this process. See Appendix B.8 for the codebook that we developed through our analysis.
@@ 4 Results
We compared LearnLM against contemporaneous flagship offerings (as of 2024-10-01), in particular GPT-4o [^4] GPT-4o version 2024-08-06, https://platform.openai.com/docs/models/gpt-4o., Claude 3.5 Sonnet [^5] Claude 3.5 Sonnet version 2024-06-20, https://docs.anthropic.com/en/docs/about-claude/models., along with Gemini 1.5 Pro [^6] Gemini 1.5 Pro-002 from 2024-09-24, https://cloud.google.com/vertex-ai/generative-ai/docs/learn/model-versions., which we adapted to train LearnLM. Since our evaluation process began, all these models have been updated and improved and new versions have been released. Therefore, our results only reflect a reasonably fair comparison at a specific moment; still, we hope that our continued investment in education maintains or increases relative preference for our models.
In total, we collected a set of $2360$ conversations, consisting of 58 4595845958\,45958 459 total learner and model messages. We collected 10 1921019210\,19210 192 expert assessments of those conversations, with an average of three experts reviewing each pair of conversations. Figure 3 shows that the systems we evaluate demonstrate notably different response length distributions across the collected conversations, including between Gemini 1.5 Pro and LearnLM, but there is no clear relationship between length and perceived quality as seen in other model comparisons [13].
@@
Figure 4: Pedagogy experts’ preferences over LearnLM and other contemporaneous systems (Claude 3.5, GPT-4o, and Gemini 1.5 Pro). The scatterplots represent the underlying distribution of seven-point preference ratings. Given the large number of ratings we collected, these scatterplots proportionally downsample to 500 ratings per measure, color-coded based on the preference scale (dark purple corresponds to strong preference for LearnLM), and randomly positioned around each integer rating for readability. The red points and error bars indicate the estimated mean and its 95% credible interval for each measure. These means are also shown in Figure 1.
We review evaluation results as follows: first, comparative preference ratings between systems from evaluation stage 3, second, non-comparative ratings from evaluation stage 3, third, non-comparative ratings after role-playing learners in evaluation stage 2, and fourth, analysis of open-ended feedback from evaluation stage 2.
@@
First, comparative preference ratings (Figure 4) reveal a strong preference toward LearnLM over GPT-4o for all five comparative assessment categories. Experts expressed the strongest preference for LearnLM in overall pedagogy (“Which tutor demonstrated better tutoring?”). They also communicated similar but smaller preferences toward LearnLM over Claude 3.5 and Gemini 1.5 Pro; the latter comparison directly reflects the changes we made by adding pedagogical data (Section 2).
@@
Figure 5: Evaluation of systems on each category of our pedagogy rubric from a 7-point Likert scale ("Strongly disagree" to "Strongly agree"). Error bars reflect 95% credible intervals from the posterior distrubtion for the mean.
@@
Figure 6: Impressions shared by the pedagogy experts role-playing as learners in our pedagogical scenarios. Error bars reflect 95% credible intervals from the posterior distribution for the mean. The rating scales for impression questions (left) were 5-point extent scales (“Not at all” to “Extremely”), and 7-point Likert scales (“Strongly disagree” to “Strongly agree”) for experience questions (right).
@@
Second, Figure 5 shows the mean performance of each model on our pedagogy rubric. Experts evaluated individual pedagogy qualities on a seven-point scale. On average, each system received a positive assessment across every rubric category from this review. LearnLM was rated highest across all rubric categories, and across almost all $29$ rubric questions, standing out on inspiring active learning, deepening metacognition, and stimulating curiosity.
@@
Third, Figure 6 depicts the degree to which each system increased participants’ interest in the tutoring topic, participants’ willingness to use the model in the future [14], and their perceptions of the competence and warmth of the model [15, 16]. Our participants reported relatively similar experiences with LearnLM, Gemini 1.5 Pro, and Claude 3.5. In contrast, participants indicated weaker experiences with GPT-4o in terms of stimulating their interest, its perceived warmth, and its perceived usefulness. Of course, these ratings come from experts role-playing learners, but they give some early indication about the user experience of these systems in the scenario settings.
@@
Fourth, we randomly subsampled 203 explanations ($20\%$ of the 1024 explanations that we collected) for thematic analysis of role-played learner preferences (see Table 1 for more details, including several example response excerpts per theme). Overall, the themes that emerged most consistently in our subsample were is_engaging (appearing in $72$ of the subsampled explanations), conversation_style ($67$ explanations), and gives_away_answers ($50$ explanations).
@@
When participants reported preferring LearnLM over the other model, their explanation was more likely to contain the themes keeps_on_topic, challenges_learner, and gives_away_answers. When participants preferred other models to LearnLM, their explanation was more likely to touch on the themes clarity, info_amount, and conversation_style. The experts role-playing as learners tended to see LearnLM as better at remaining on topic and guiding learners to a robust understanding of concepts, rather than simply giving away answers. On the other hand, these experts occasionally found LearnLM to be less suitable in terms of information delivery or conversation style.
Theme Appearances when participants preferred LearnLM ($n=94$) Appearances when participants preferred other models ($n=80$) Example responses
keeps_on_topic $20\leavevmode\nobreak\ (\mathbf{21.2\%})$ $8\leavevmode\nobreak\ (10\%)$ “[LearnLM] didn’t let me get away with distractions”
@@
“[The other tutor] also did a much better job of getting me back on task”
challenges_learner $31\leavevmode\nobreak\ (\mathbf{33.0\%})$ $13\leavevmode\nobreak\ (16.3\%)$ “obviously [LearnLM] was better […] [the other tutor] clearly wasn’t pushing me to do well”
@@
“[The other tutor] asked interesting questions that made me think deeper”
gives_away_answers $32\leavevmode\nobreak\ (\mathbf{34.0\%})$ $15\leavevmode\nobreak\ (18.8\%)$ “[LearnLM] really engaged me in the steps to answer the question whereas [the other tutor] just gave me the answer”
@@
“[LearnLM] was too reticent to help by giving answers when it was clear the student needed it”
clarity $15\leavevmode\nobreak\ (16.0\%)$ $16\leavevmode\nobreak\ (\mathbf{20.0\%})$ “The structure of the support [for the other tutor] was a bit clearer for the student to follow”
@@
“I just thought the answers [for LearnLM] were more clear”
info_amount $19\leavevmode\nobreak\ (20.2\%)$ $20\leavevmode\nobreak\ (\mathbf{25.0\%})$ “[The other tutor] was […] more succinct”
@@
“[LearnLM] did a better job of breaking this "complex" topic into more digestible bites”
conversation_style $30\leavevmode\nobreak\ (31.9\%)$ $29\leavevmode\nobreak\ (\mathbf{36.3\%})$ “I […] felt that [LearnLM] was a bit patronizing”
@@
“[LearnLM] was warmer and more encouraging”
Table 1: Themes which were more likely to appear in “learner” explanations of preferences favoring LearnLM (top three rows), or favoring other models (bottom three rows). This table displays themes (i) referenced by at least $10\%$ of all sampled preference explanations, and (ii) showing an extreme ratio of occurrence between explanations favoring LearnLM and explanations favoring other models.

### v3

@@ Abstract
Today’s generative AI systems are tuned to present information by default, rather than engage users in service of learning as a human tutor would. To address the wide range of potential education use cases for these systems, we reframe the challenge of injecting pedagogical behavior as one of pedagogical instruction following, where training and evaluation examples include system-level instructions describing the specific pedagogy attributes present or desired in subsequent model turns. This framing avoids committing our models to any particular definition of pedagogy, and instead allows teachers or developers to specify desired model behavior. It also clears a path to improving Gemini models for learning—by enabling the addition of our pedagogical data to post-training mixtures—alongside their rapidly expanding set of capabilities. Both represent important changes from our initial tech report [1]. We show how training with pedagogical instruction following produces a LearnLM model (available on Google AI Studio) that experts substantially prefer across a diverse set of learning scenarios, with average preference strengths of +31% over GPT-4o, +11% over Claude 3.5 Sonnet, and +13% over the Gemini 1.5 Pro model on which LearnLM was based.
@@ 1 Introduction
Figure 1: An overview of our three-stage expert evaluation pipeline and our results for comparing LearnLM with other systems. (1) We developed learning scenarios that allow expert participants to role-play specific learners interacting with pairs of AI tutors. (2) Grounding material (e.g. an essay, homework problem, diagram, etc.) and System Instructions specific to each scenario are passed as context to each model. We pass the resulting conversation pairs to pedagogy experts (3) who review each model on its own, as well as their comparative performance. We aggregate these comparative assessments (on a seven-point response scale) (4) to evaluate overall preference for LearnLM over GPT-4o, Claude 3.5 Sonnet, and Gemini 1.5 Pro. See Section 4 for more detailed results.
Using the updated methodology, we trained a new version of LearnLM, based on Gemini 1.5 Pro [^2] Specifically gemini-1.5-pro-002 (release notes). [3]. In our evaluations against contemporaneous flagship models, each representing a company’s premier offering as of 2024-10-01, educators and pedagogy experts preferred this version of LearnLM with an average preference strength of $31\%$ over GPT-4o, $11\%$ over Claude 3.5 Sonnet, and $13\%$ over the original Gemini 1.5 Pro (see Figure 1). LearnLM is available as an experimental model on Google AI Studio along with documentation of example use cases and suggested prompts. We welcome any feedback on LearnLM to help inform our future research and priorities. As we improve LearnLM for teaching and learning, we are also working to bring these advances into Gemini models, so any developers using Gemini can benefit from the improvements made via LearnLM research.
Section 2 describes how we trained LearnLM for pedagogical instruction following and Section 3 explains how we updated our scenario-based evaluation design accordingly. Section 4 shows a detailed analysis of results comparing LearnLM with other premier model offerings. Finally, Section 5 outlines some future work, especially with regards to continued evaluation. In addition to the broad range of core academic subjects that we use for training and evaluation, we include a feasibility study on medical education subjects in Appendix C.
@@ 2.3 Benefits of co-training
3 Expert Evaluation Design
In our initial tech report, we discussed a taxonomy of pedagogy evaluation designs and reported results of four human evaluations with different methodologies (Sections 4 and 5 in Jurenka et al. [1]). Within this taxonomy, here we focus on scenario-guided, conversation-level pedagogy evaluations and side-by-side comparisons. For this new set of evaluations, we improved the clarity and coverage of our learning scenarios, added System Instructions specific to each scenario, and updated our pedagogy rubric and questions. Guiding participant conversations with scenarios is especially important in multi-turn settings [10]. Without scenarios, the unconstrained nature of human-AI interactions frequently leads to meandering conversations, offering a poor basis for comparison. In contrast, scenario-based approaches support relatively repeatable, controlled comparisons of the capabilities of different conversational AI systems. Scenario frameworks also help with evaluation coverage, ensuring that we test a diverse range of use cases.
Our evaluation process takes place in three stages, depicted above in Figure 1. First, we identified an ecologically representative distribution of learning use cases and created a bank of 49 evaluation scenarios (Section 3.1). Second, these scenarios grounded interactions between AI systems and a pool of $N=186$ educators and pedagogy experts role-playing as learners across learning goals, subjects, learning materials, and learner personas (Section 3.2). Third, to assess the quality of pedagogy in these interactions, we separately recruited a pool of $N=248$ educators and pedagogy experts to review the performance of the systems (Section 3.3). This process produced ample quantitative and qualitative data to help us understand the systems’ capabilities and behavior (Section 3.4).
We are committed to following best practices in research ethics, including by communicating transparently about our research aims, collecting informed consent, and compensating fairly for participation [11]. Our protocol underwent independent ethical review, with a favourable opinion from the Human Behavioural Research Ethics Committee at Google DeepMind (#23 011).
@@ 3.1 Scenario design
We next collaboratively and iteratively populated our bank of scenarios. Members of our team—including two with years of professional experience educating students and training teachers—independently drafted scenarios, leveraging the template and guiding questions from Phase 2. We collectively reviewed the scenario drafts, assessing each for clarity, completeness, correctness, and relevance to our pedagogical principles and the use cases defined in Phase 1. We weighted the overall distribution of scenarios across different learning goals, personas, and subject areas, flagging any gaps for further development.
This process resulted in a diverse bank of 49 scenarios across core academic subjects (for examples, see Appendix B.3). In addition to building this foundational bank, we later validated the robustness and replicability of this procedure by running a feasibility study in a specialized educational domain (specifically, medical education; see Appendix C).
@@ 3.2 Conversation collection
Every session of conversation collection began with a short training on role-playing the scenarios (see Figure 2, Step 1). After passing a quiz at the end of the training, participants selected a scenario to enact (see Figure 2, Step 2). Conversation collection proceeded in pairs, such that the same participant enacted a scenario first with one AI system, and then another. Each pair included LearnLM and a comparison system. We randomized the order of the systems for each conversation pair and did not label the systems for participants. Within each pair of conversations, the models received the same System Instructions, grounding material, and initial learner queries as context, as specified by our scenarios (see Figure 2, Step 3). We formatted all inputs identically, except for some small specification differences mandated by the system APIs.
As specified by our template, each scenario included an initial query for the learner that was automatically sent on behalf of the participant to begin the conversation. After the AI system responded to that query, the participant continued the interaction, guided by the information provided in the scenario. We required participants to continue for a minimum of 10 conversational turns (thus, a minimum of five learner and five tutor turns) before they could end the interaction.
After ending each conversation (see Figure 2, Step 4), participants filled out a brief questionnaire to share their experience interacting with the system (see Figure 2, Step 5 & Appendix B.4). Additionally, after each pair of conversations, participants completed another questionnaire focused on comparing their impressions of the two systems (see Figure 2, Step 6 & Appendix B.5). Participants could then either select a new scenario to begin two additional conversations or end the session.
@@ 3.4 Analysis
Our study design involves repeated measurements from our participants. That is, each participant role-playing as a learner interacted with each system multiple times, and each expert assessed each system multiple times. To account for this non-independence and avoid artificially inflating our certainty in our estimates, we analyze our data with hierarchical regressions [12]. Appendix B.8 describes our statistical methods in more detail.
In addition, we conducted qualitative analysis of the open-ended comments and feedback collected from our experts after role-playing each scenario with two systems (Stage 2) [^3] For our qualitative analysis, we prioritized the feedback from Stage 2 for its unique insights into the firsthand experience of learning with each tutor.. In particular, we identified and then refined general themes related to the learner-system interactions from participants’ free-form responses. We then coded individual responses for the presence or absence of each theme. To avoid biasing our annotations, we censored the identities of the systems during this process. Appendix B.9 presents the codebook that we developed through our analysis.
@@ 4 Results
We compared LearnLM against contemporaneous flagship offerings (as of 2024-10-01), in particular GPT-4o [^4] GPT-4o version 2024-08-06, https://platform.openai.com/docs/models/gpt-4o., Claude 3.5 Sonnet [^5] Claude 3.5 Sonnet version 2024-06-20, https://docs.anthropic.com/en/docs/about-claude/models., and Gemini 1.5 Pro [^6] Gemini 1.5 Pro-002 from 2024-09-24, https://cloud.google.com/vertex-ai/generative-ai/docs/learn/model-versions.. Since we undertook this specific evaluation, each of these models has been updated, with new versions released. Our results should therefore be understood as a point-in-time comparison, evaluating the effectiveness of our approach and establishing a baseline for our continued investment in education.
In total, we collected a set of $2360$ conversations, consisting of 58 45958\,459 total learner and model messages. We collected 10 19210\,192 expert assessments of those conversations, with an average of three experts reviewing each pair of conversations. Figure 3 shows that the systems we evaluate demonstrate notably different response length distributions across the collected conversations, including between Gemini 1.5 Pro and LearnLM. On an aggregate level, we observe no clear relationship between length and perceived quality (cf. [13]).
@@
Figure 4: Pedagogy experts’ preferences over LearnLM and other contemporaneous systems (Claude 3.5 Sonnet, GPT-4o, and Gemini 1.5 Pro). The scatterplots represent the underlying distribution of seven-point preference ratings. Given the large number of ratings we collected, these scatterplots proportionally downsample to 500 ratings per measure, color-coded based on the preference scale (dark purple corresponds to strong preference for LearnLM), and randomly jittered around each scale value for readability. The red points and error bars indicate the estimated mean and its 95% credible interval for each measure.
We begin our analysis by examining the pedagogical assessments and preference ratings from our expert evaluators. We then explore the firsthand feedback from participants who role-played as learners interacting with the models. That is, we present our pedagogical findings from Stage 3 before backtracking to the interaction data gathered from Stage 2. Several clear patterns emerge from this analysis.
@@
First, comparative preference ratings (Figure 4) reveal a strong preference toward LearnLM over GPT-4o for all five comparative assessment categories. Experts expressed the strongest preference for LearnLM in overall pedagogy (“Which tutor demonstrated better tutoring?”). They also communicated similar but smaller preferences toward LearnLM over Claude 3.5 Sonnet and Gemini 1.5 Pro. Because we adapted Gemini 1.5 Pro to train LearnLM, the comparison between those two models directly reflects the changes that result from adding pedagogical data (see Section 2).
@@
Figure 5: Evaluation of systems in each category of our pedagogy rubric on a seven-point response scale (“Strongly disagree” to “Strongly agree”). Error bars reflect 95% credible intervals from the posterior distrubtion for the mean.
@@
Figure 6: Impressions shared by the pedagogy experts role-playing as learners in our pedagogical scenarios. Error bars reflect 95% credible intervals from the posterior distribution for the mean. Participants responded to the impression items on a five-point response scale (“Not at all” to “Extremely”) and to the experience items on a seven-point response scale (“Strongly disagree” to “Strongly agree”).
@@
Second, Figure 5 shows the mean performance of each model on our pedagogy rubric. Experts evaluated individual pedagogy qualities on a seven-point scale. On average, each system received a positive assessment across every rubric category from this review. Experts collectively assigned the highest marks to LearnLM across all rubric categories, and across almost all $29$ rubric questions, with particularly large leads on inspiring active learning, deepening metacognition, and stimulating curiosity.
@@
Third, Figure 6 depicts the degree to which each system increased participants’ interest in the tutoring topic, participants’ willingness to use the model in the future [14], and their perceptions of the competence and warmth of the model [15, 16]. Our participants reported relatively similar experiences with LearnLM, Gemini 1.5 Pro, and Claude 3.5 Sonnet. In contrast, participants indicated weaker experiences with GPT-4o in terms of its effects on their interest, its perceived warmth, and its perceived usefulness. While role-playing experts represent imperfect proxies for students, their impressions help provide preliminary insight into the user experience of AI tutoring interactions.
@@
Fourth, we randomly subsampled 203 explanations (approximately $20\%$ of the 1024 explanations that we collected) for thematic analysis of role-played learner preferences (see Table 1 for more details, including several example excerpts per theme). Overall, the themes that emerged most consistently in our subsample were is_engaging (appearing in $72$ of the subsampled explanations), conversation_style ($67$ explanations), and gives_away_answers ($50$ explanations).
@@
When participants reported preferring LearnLM over the other model, their explanation was more likely to contain the themes keeps_on_topic, challenges_learner, and gives_away_answers. In contrast, participants who preferred other models to LearnLM tended to touch on the themes of clarity, info_amount, and conversation_style in their explanations. Overall, our community of experts tended to see LearnLM as better at remaining on topic and guiding learners to a robust understanding of concepts, rather than simply giving away answers. On the other hand, these experts occasionally found LearnLM to be less suitable in terms of information delivery or conversation style.
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
Table 1: Themes that were more likely to appear in “learner” explanations of preferences favoring LearnLM (top three rows), or favoring other models (bottom three rows). This table displays themes (i) referenced by at least $10\%$ of all sampled preference explanations, and (ii) showing an extreme ratio of occurrence between explanations favoring LearnLM and explanations favoring other models.